40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10731     Entry time: Fri Nov 21 13:58:51 2014
Author: ericq 
Type: Update 
Category: Optical Levers 
Subject: HeNe RIN test 

 Steve had me measure the RIN of a JDSU HeNe laser. I used a PDA520, and measured the RIN after the laser had been running for about an hour, which let the laser "settle" (I saw the low frequency RIN fall after this period).

Here's the plot and zipped data.

Steve: brand new laser with JDSU 1201 PS

RIN_P893519.pdf

Attachment 1: 2014-11-21_HeNeRIN.zip  186 kB
ELOG V3.1.3-